# SonicVisionLM: Playing Sound with Vision Language Models

### v2

@@ Abstract
There has been a growing interest in the task of generating sound for silent videos, primarily because of its practicality in streamlining video post-production. However, existing methods for video-sound generation attempt to directly create sound from visual representations, which can be challenging due to the difficulty of aligning visual representations with audio representations. In this paper, we present SonicVisionLM, a novel framework aimed at generating a wide range of sound effects by leveraging vision language models. Instead of generating audio directly from video, we use the capabilities of powerful vision language models (VLMs). When provided with a silent video, our approach first identifies events within the video using a VLM to suggest possible sounds that match the video content. This shift in approach transforms the challenging task of aligning image and audio into more well-studied sub-problems of aligning image-to-text and text-to-audio through the popular diffusion models. To improve the quality of audio recommendations with LLMs, we have collected an extensive dataset that maps text descriptions to specific sound effects and developed temporally controlled audio adapters. Our approach surpasses current state-of-the-art methods for converting video to audio, resulting in enhanced synchronization with the visuals and improved alignment between audio and video components. Project page: https://yusiissy.github.io/SonicVisionLM.github.io/
Figure 1: A model implements the automatic detection of in-picture voice generation and accepts the user’s editing of this paper and time in the off-screen section.On-screen sound means sound originate from visual information.Off-screen sound not directly happened on the screen.
Figure 2: SonicVisionLM’s framework. SonicVisionLM presents a composite framework designed for the automatic recognition of on-screen sounds coupled with a user-interactive module for off-screen sounds editing. The blue dashed box and arrows in the figure represent the visual automation workflow: First, a silent video goes into the system where it figures out what’s happening (text)and when(time). Then, this info is used to make the sounds that match what’s on screen. The purple dotted box and arrows show how users can add their own changes to create off-screen sounds.
@@ 1 Introduction
The sound effects artists work with different types of sounds, some you could see on-screen (like footsteps or a car passing) you couldn’t see but hear (like background noises that make the video feel more real and add to the story). On-screen sounds match what’s happening in the video, while off-screen sounds help set the mood and give extra information. Making sounds for videos is a key part of video production, however, it could take lots of labor for the artists. As a result, the video-sound generation task has gained significant attention.
Although the recent approaches have made great efforts, the video-sound generation task is still challenging. For on-screen sounds, achieving semantic relevance and maintaining temporal synchronization continues to be a complex issue. For off-screen sounds, it is hard to edit. Current methods [31, 13, 24] primarily focus on the visual content to generate the corresponding sound, a subset of these methods [8, 7] takes into account the aspect of editability. However, the alignment between video features and audio features is quite difficult, leading to deficiencies including 1) incorrect sound meanings, and mismatched timing, 2) monotonous sound effects, and lack complex scenarios. Both lead to unsatisfactory results in video-sound generation task.
To solve the above deficiency, we propose a novel framework, named SonicVisionLM, as shown in Fig.1. SonicVisionLM is proposed by introducing three key components, including video-to-text, text-based interaction, and text-to-audio generation. First, the video-to-text component focuses on generating sound effects for on-screen events. In this step, VLM is employed to identify pertinent sound descriptions from the input silent video. Following this, a temporal detection network is trained to extract specific text-based interact temporal information from the video. A key innovation within this framework is the design of a time-conditioned embedding, which is utilized to guide an audio adapter. After that, the text-based interaction component allow users to change the text and timestamps from a previous video-to-text component, or to input new corresponding text-timestamp pairs for personalized sound design. Finally, the text-to-audio generation component accept the text and timestamp conditions, and then input them into the LDM and adapter to generate diverse, time-synchronized, and controllable sounds. Simultaneously, We collect a text-to-single-sound dataset, named CondPromptBank, for sound effects caption and timing cues, comprising over ten thousand data points, covering 23 different categories, CondPromptBank can generate high-quality and divers sound. The main contributions of this work are:
  • We propose a novel framework called SonicVisionLM and collect a dataset CondPromptBank specifically for training a time-controllable adapter. This ensures the generated sound aligns perfectly with our text input and maintains precise timing control.
@@
  • The proposed framework achieves state-of-the-art result on both conditional and unconditional video-sound generation tasks.The conditional task can be noticeably enhanced in all metrics. (IoU: 22.4$\rightarrow$39.7)
@@ 2 Related Work
We roughly divide audio generation tasks into two categories. The domain of Text-to-Audio generation encompasses various subfields, including Text-to-Speech (TTS) and Text-to-Music (TTM). State-of-the-art TTS models like FastSpeech2 [28] and NaturalSpeech [32] have made significant progress, producing speech quality nearly indistinguishable from human speech. In the TTM domain, MusicLM [1], Noise2Music [11], MusicGen [6], and MeLoDy [20] are aimed to generate music segments from textual descriptions, bringing significant innovation to the field of music composition and synthesis. Models like AudioGen [18], AudioLDM [23], Tango[9], and Make-an-Audio [12] focus on universal audio generation modeling. AudioGen [18]treats audio generation as a conditional language modeling task, while the other three models employ latent diffusion methods to accomplish sound generation. However, due to the comprehensive large-scale datasets, existing models lack the direct capability to generate sound effects for videos. Therefore, we created a text-to-single-sound dataset to help the model produce high-standard sound effects. In Video-to-Sound Generation task, SpecVQGAN [13]employs a Transformer-based autoregressive model, utilizing extracted ResNet50 or RGB+Flow features for sound generation. Im2Wav [31]uses a dual-transformer model conditioned on CLIP features for sound generation. CondFoleyGen [8]and VARIETYSOUND [7]have introduced tasks for controllable timbre generation. Diff-foley[24] uses contrastive audio-visual pretraining to align audio and visual features However, the results generated by these methods have poor audio-visual synchronization, obvious noises, and uneditable. Unlike the previously mentioned works, our model not only audio-visual synchronization and diversity, but also allows for personalized edits from users. Our model provides a more comprehensive sound solution for video production.
@@
Diffusion Model. have been explored for both mel-spectrogram generation [27, 4], and waveform generation [19, 21, 3]. However, the iterative generation process can be slow for high-dimensional data. AudioLDM[23], Make-An-Audio[12], and Tango[9], which trained in a continuous latent space. However, the control of LDM in audio generation tasks still falls short of achieving satisfactory results. In this paper, we attempt to control the timing of sound generation and ensure audio-visual synchronization.
@@
Vision Language Models. Recently, GPT-4 has showcased advanced multi-modal abilities and inspired vision-language LLMs. Vision-LLM [35] and LLaVA[33] are aligning image inputs with large language models Vicuna[5] exhibit similar multi-modal capabilities. Recent developments in this field include MiniGPT-v2[2]. Kosmos-2 [26] demonstrating the ability of multi-model LLMs to perform visual grounding. In this paper, we first introduce the VLM to the audio generation task.
@@ 3.1 Overview
In this section, we introduce the framework of SonicVisionLM, as shown in Fig. 2. Before delving into the specific design details, we first provide a brief overview of the preliminary knowledge (Section 3.1). Then, we introduce the Visual-to-Audio Event Detection Module (Section 3.2), which obtains textual descriptions of on-screen sounds through VLMs. Subsequently, we present the Sound Event Timestamp Detection Module (Section 3.3), designed to accurately detect the timing cues through network architecture. Finally, we introduce the proposed time-controllable adapter as an extension of the audio diffusion model (Section 3.4), enabling the generation of multiple sounds that are semantically coherent and temporally aligned.
@@ 3.2 Preliminaries
Audio Diffusion Model. The text-prompt encoder encodes the input description $\tau\in R^{L\times d_{t}}$ of the sound, where $L$ is the token count and $d_{t}$ and is token-embedding size.
The latent diffusion model (LDM) to construct the audio prior $z_{0}$ with the guidance of text encoding $\tau$. This essentially reduces to approximating the true prior $q\left(z_{0}\mid\tau\right)$ with parameterized $p\left(z_{0}\mid\tau\right)$. LDM can achieve the above through forward and reverse diffusion processes. The forward diffusion is a Markov chain of Gaussian distributions with scheduled noise parameters $10<\beta_{1}<\beta_{2}<\textperiodcentered\textperiodcentered\textperiodcentered<\beta_{N}<$ 1 to sample noisier versions of $z_{0}$, where $N$ is the number of forward diffusion steps, $\alpha_{n}=1-\beta_{n}$, and $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{n}$ :
@@
where the noise term $\epsilon\sim\mathcal{N}(0,I)$. The final step of the forward process yields $z_{N}\sim\mathcal{N}(0,I)$.
The reverse process denoises and reconstructs $z_{0}$ through text-guided noise estimation $(\hat{\epsilon}_{\theta})$ using loss:
@@
After LDM is trained, we generate audio latent by sampling through the reverse process with $z_{N}\sim\mathcal{N}(0,I)$, conditioned on the given textual representation $\tau$, with the following reverse dynamics:
@@
Thereafter the decoder of audio VAE [17] constructs a mel-spectrogram feature from the latent audio representation $\hat{z}_{0}$. This mel-spectrogram feature is convey to a vocoder to generate the final audio.
@@ 3.3 Visual-to-Audio Event Understanding Module
The diversity of sound is influenced by various factors, including the source of the sound, actions, the environment, and more. Common methods of naming sound effects are also based on these factors. Inspired by the widespread use of visual language models[22, 36, 38], we chose miniGPT-v2[2]to process visual information and generate descriptions of sounds. Due to miniGPT-v2 is originally designed for single-image understanding, recognizing the limitation of single-image in conveying dynamic information, We follow the LLaMA-2[33] conversation template design and adapt it for the multi-modal instructional template. we made adjustments to accept the four random video frames feature and incorporated temporal cues in the instructions. The adapted template is defined as follows:
$First,\ \ \ \ \ \ \ \ <Img><ImageFeature><Img>.\\ \ Then,\ \ \ \ \ \ \ \ <Img><ImageFeature><Img>.\\ \ Afterthat,<Img><ImageFeature><Img>.\\ \ Finally,\ \ \ \ \ <Img><ImageFeature><Img>.\\ \left[Task\ \ Identifier\right]\ \ \ Instruction$
@@ 3.4 Sound Event Timestamp Detection Module
We use the sound event timestamp detection module is used to detect the timestamp of the sound event in the video inspired by the CondFoleyGen[8]. Instead of using a hand-crafted approach to transfer sounds from the conditional example, We train a ResNet (2+1)-D18[34] visual network to capture the timestamp of on-screen sound events. We pre-trained this network using a dataset comprising videos and sound timing data. The dataset generates the ground truth timestamp with the timing label that is detected by aligning it to the closest frame. We use a binary cross-entropy loss, penalizing time predictions that occur at incorrect times. Since each input video can have more than one sound event, we set the weight of each onset in the loss from the video equally according to the total number of onsets in the input so that the weight sums to one. Subsequently, we input the silent video frame sequence $V_{f}$ into the detection network, and the output is a one-hot vector $V_{ct}$ of predictions using a fully connected layer after pooling (one prediction per input frame). This vector serves as input for the subsequent Time-controllable Latent Diffusion Model. The whole process and network structure are shown in Fig.3.
Figure 3: Sound Event Timestamp Detection Module.The network organizes the video’s features to create a special heat vector. This vector is as long as the number of frames in the video. The white parts in it show where there’s sound, and the black parts show where there’s no sound.
@@ 3.5 Time-controllable Latent Diffusion Model
In our experimentation, we observed that the results generated were semantically inaccurate and temporally unsynchronized. It would happen when such methods employed audiovisual datasets for training end-to-end models. This problem hinders the widespread application of this task in reality. We posit that to solve effectively aligning visual features with corresponding audio features could solve this problem. Therefore, we propose using timestamps to control the text-to-sound generation model, which improves sound quality, diversity, and synchronization.
In the visual domain, many works are based on the ControlNet architecture, which can finer control the generation of images or videos by manipulating the input image conditions of neural network blocks. However, unlike the visual domain’s inherent intuitiveness, audio features exhibit temporal continuity and are inherently more complex and abstract. Consequently, the selection of appropriate audio features to guide the generation process presents a significant challenge. In this paper, we propose a embedding called audio time-condition embedding $A_{ct}$, which involves the following steps: Firstly, we use function $F(\cdot)$ extract the start $x_{start}$ and end points $x_{end}$ of sound segments from the audio $X_{a}$ and expand them into one-hot vectors $V_{ct}$of the same length as the audio, where these one-hot vectors indicate the presence of sound in each audio frame.
$x_{start},x_{end}=F\left(a_{c}\right)$ (7)
$V_{ct}=\begin{cases}1\ \ \ ,if\ \ t\ \ in\ \ \ [x_{start},x_{end}]\\ 0\ \ \ ,else\end{cases}$ (8)
Subsequently, based on these one-hot vectors $V_{c}t$, we fill all channels of the mel-spectrogram with maximum values at the corresponding time frames Finally, we obtain mel feature embeddings $A_{ct}$ through a VAE encoder $E_{a}$.
$a_{ct}=E_{a}\left(\operatorname{max}\left(\operatorname{mel}\left(a_{c}\right),V_{ct}\right)\right)$ (9)
we have taken inspiration from ControlNet[39], designed a network architecture,called Time-controllable Adapter. Then, the audio time-condition embedding $A_{ct}$ is with the text embedding $\tau$and target audio embedding, fed into the neural network block to train the time-controllable adapter together. For brevity, we denote the encoder as $F$ , the middle block as $M$, and the decoder as $G$, with $f_{i}$ and $g_{i}$denoting the output of the $i-th$ block in the encoder and decoder, and $m$ denoting the output of the middle block, respectively. Specifically, we fix the weights of Tango and copy the structures and weights of the encoder and middle block, designated as $F^{\prime}$ and $M^{\prime}$respectively. Thereafter, we incorporate the information from the time-controllable adapter during the decoding process. The output of the time-controllable adapter is added to the output of the corresponding block of the decoder. To achieve it, we ensure that all other elements are kept frozen while modifying the input of the $i-th$ block of the decoder as:
@@
where $zero$ represents one zero convolutional layer whose weights increase from zero to gradually integrate control information into the main diffusion model. This approach was adopted to preserve the audio generation capabilities of the large-scale model, which was trained on billions of audio-text pairs. Simultaneously, it enabled the model to comprehend the guidance provided by the time control embedding, resulting in temporally controllable outcomes.
[…]
$*{}^{*}$[8] 15.5 21.6 67.0 35.5 23.3
CondFoleyGen(codebook) $*{}^{*}$[8] 16.3 19.2 68.6 37.4 22.4
Ours-small 21.5 19.4 77.03 27.8 35.6
Ours-full 46.7 27.6 78.1 43.8 39.7
Table 1: Conditional Generation Task Qualitative Results. CLAP-top1 is a metric that evaluates the model’s conditional controllability. It calculates the percentage of times the sound samples generated by the four models rank in the top 1 according to CLAP rankings, divided by the total number of samples. Onset Acc and Time Acc are both metrics based on the number of sound occurrences, with Onset Acc focusing on the onset count, and Time Acc on the count of time intervals. We measure the average accuracy of predicting onsets within 0.1 seconds of ground truth to assess the timing of generated onsets. IoU is calculated by computing the intersection and union of these vectors. These metrics collectively allow us to comprehensively evaluate the accuracy of the generated sounds in terms of both timing and content. $*{}^{*}$ ”denotes the data sourced from the official code and is based on experiments conducted with our local configurations. underline denotes the worst performance. boldface denotes the best performance. The five metrics all measured in percentage.

### v3

@@ Abstract
There has been a growing interest in the task of generating sound for silent videos, primarily because of its practicality in streamlining video post-production. However, existing methods for video-sound generation attempt to directly create sound from visual representations, which can be challenging due to the difficulty of aligning visual representations with audio representations. In this paper, we present SonicVisionLM, a novel framework aimed at generating a wide range of sound effects by leveraging vision-language models(VLMs). Instead of generating audio directly from video, we use the capabilities of powerful VLMs. When provided with a silent video, our approach first identifies events within the video using a VLM to suggest possible sounds that match the video content. This shift in approach transforms the challenging task of aligning image and audio into more well-studied sub-problems of aligning image-to-text and text-to-audio through the popular diffusion models. To improve the quality of audio recommendations with LLMs, we have collected an extensive dataset that maps text descriptions to specific sound effects and developed a time-controlled audio adapter. Our approach surpasses current state-of-the-art methods for converting video to audio, enhancing synchronization with the visuals, and improving alignment between audio and video components. Project page: https://yusiissy.github.io/SonicVisionLM.github.io/
Figure 1: A model implements the automatic detection of on-screen sound generation and accepts the user’s editing of text and time in the off-screen section. On-screen sound refers to audio that originates from visible actions within the video frame. Off-screen sound is not directly observable on the screen.
Figure 2: SonicVisionLM’s framework. SonicVisionLM presents a composite framework designed to automatically recognize on-screen sounds coupled with a user-interactive module for editing off-screen sounds. The blue dashed box and arrows in the figure represent the visual automation workflow: First, a silent video goes into the system to determine the occurring events (text) and their timing (time). Then, this information conditions the generation of sounds matching the screen. The purple dotted box and arrows show how users can modify or add off-screen sounds.
@@ 1 Introduction
The sound effects artists work with various types of sounds, including those visible on-screen (like footsteps or a car passing) and those audible but not visible (like background noises and heartbeats can enhance the video’s authenticity and narrative). On-screen sounds match what is happening in the video, while off-screen sounds establish the ambience and provide additional information. Creating soundtracks for videos is a vital aspect of video production, but it can be labour-intensive for artists. Therefore, the video-sound generation task has gained notable attention.
Although the recent approaches have made great efforts, the video-sound generation task is still challenging. For on-screen sounds, achieving semantic relevance and maintaining temporal synchronization continues to be a complex issue. It is hard to edit off-screen sounds. Current methods [32, 14, 25] primarily focus on the visual content to generate the corresponding sound, a subset of these methods [9, 7] considers editability. Nonetheless, the alignment between video and audio features is tricky, leading to deficiencies including 1) incorrect sound meanings and mismatched timing, 2) monotonous sound effects, and a lack of complex scenarios. Both lead to unsatisfactory results in the video-sound generation task.
We propose a novel framework named SonicVisionLM to solve the above deficiency, as shown in Fig. 1. SonicVisionLM is proposed by introducing three key components: video-to-text, text-based interaction, and text-to-audio generation. First, the video-to-text component focuses on generating sound effects for on-screen events. This step uses a VLM to identify appropriate sound descriptions from the input silent video. Following this, a timestamp detection network is trained to extract specific temporal information from the video. A key innovation within this framework is the design of a time-conditioned embedding, which is utilized to guide an audio adapter. After that, the text-based interaction component allows users to change the text and timestamps from a previous video-to-text component or to input new corresponding text-timestamp pairs for personalized sound design. Finally, the text-to-audio generation component accepts the text and timestamp conditions and inputs them into the LDM and adapter to generate diverse, time-synchronized, and controllable sounds. Simultaneously, We collect a text-to-single-sound dataset, named CondPromptBank, for sound effects caption and timing cues, comprising over ten thousand data points, covering 23 categories. The main contributions of this work are:
  • We propose a novel framework called SonicVisionLM and collect a dataset CondPromptBank specifically for training a time-controllable adapter. It ensures the generated sound aligns perfectly with our text input and maintains precise timing control.
@@
  • The proposed framework achieves state-of-the-art results on conditional and unconditional video-sound generation tasks. The conditional task can be noticeably enhanced in all metrics. (IoU: 22.4$\rightarrow$39.7)
@@ 2 Related Work
Audio Generation can be broadly classified into two categories. The Text-to-Audio Generation field includes Text-to-Speech (TTS) and Text-to-Music (TTM). Leading TTS models, such as FastSpeech2 [29] and NaturalSpeech [33], now produce speech virtually indistinguishable from human speech. In TTM, MusicLM [1], Noise2Music [12], MusicGen [6] and MeLoDy [21] are aimed to generate music segments from text, bringing innovation to music composition and synthesis. Models like AudioGen [19], AudioLDM [24], Tango [10], and Make-an-Audio [13] focuses on universal audio generation modeling. AudioGen [19] treats audio generation as a conditional language modelling task, while the other three models employ latent diffusion methods to accomplish sound generation. Current methods use datasets including sound effects, voices, and music, but practical applications use these elements separately. As the textual descriptions of time are subjective, videos are more intuitive and precise, so V2A requires more accurate semantic features and time control than T2A. Therefore, we have created a text-to-single-sound dataset called CondPromptBank with detailed semantic segmentation and temporal annotations that help models produce high-quality sound effects for videos. In Video-to-Audio Generation task, SpecVQGAN [14] utilizes a Transformer-based autoregressive model, drawing on ResNet50 or RGB+Flow features to generate sound. Im2Wav [32] uses a dual-transformer model conditioned on CLIP features for sound generation. CondFoleyGen [9] and VARIETYSOUND [7] introduce tasks for controllable timbre generation. Diff-foley [25] uses contrastive audio-visual pretraining to align audio and visual features. ClipSonic [8] learns the text-audio correspondence by leveraging the audio-visual correspondences in videos and the multi-modal representation learned by pre-trained VLMs. However, the sounds generated by these methods often suffer from poor audio-visual synchronization, noticeable noises, and lack of editability. Unlike the works above, our model ensures audio-visual synchronization, enriches diversity, and supports personalized user edits. Our model provides a more comprehensive sound solution for video production.
@@
Diffusion Model has been utilized for generating both mel-spectrogram generation [28, 4], and waveform generation [20, 22, 3]. However, their iterative generation process can be slow for high-dimensional data. Models such as AudioLDM [24], Make-An-Audio [13], and Tango [10] have successfully trained diffusion models within a continuous latent space. Nevertheless, achieving satisfactory results in controlling LDM for audio generation tasks remains challenging. This paper aims to introduce time control to ensure audio-visual synchronization.
@@
Vision Language Models like ChatGPT-4, which demonstrated advanced multi-modal abilities and inspired vision-language LLMs. Vision-LLM [36] and LLaVA [34] focus on aligning image inputs with large language models Vicuna[5] exhibit similar multi-modal capabilities. Recent developments in this field include MiniGPT-v2 [2]. Kosmos-2 [27] demonstrates multi-modal LLMs’ ability to perform visual grounding. In this paper, we first introduce the VLMs to the audio generation task.
@@ 3.1 Overview
In this section, we introduce the framework of SonicVisionLM, as shown in Fig. 2. Before delving into the specific design details, we first briefly overview the preliminary knowledge (Sec. 3.2). Then, we introduce the Visual-to-Audio Event Detection Module (Sec. 3.3), which obtains textual descriptions of on-screen sounds through VLMs. Subsequently, we present the Sound Event Timestamp Detection Module (Sec. 3.4), designed to accurately detect the timing information through network architecture. Finally, we introduce the proposed time-controllable adapter as an extension of the audio diffusion model (Sec. 3.5), enabling the generation of multiple sounds that are semantically coherent and temporally aligned.
@@ 3.2 Preliminaries
Audio Diffusion Model. The text-prompt encoder encodes the input description $\tau\in R^{L\times d_{t}}$ of the sound, where $L$ is the token count and $d_{t}$ and is the token-embedding size. The latent diffusion model (LDM) is used to construct the audio prior $z_{0}$ with the guidance of text encoding $\tau$. This essentially reduces to approximating the true prior $q\left(z_{0}\mid\tau\right)$ with parameterized $p\left(z_{0}\mid\tau\right)$. LDM can achieve the above through forward and reverse diffusion processes. The forward diffusion is a Markov chain of Gaussian distributions with scheduled noise parameters $10<\beta_{1}<\beta_{2}<\cdot\cdot\cdot<\beta_{N}<$ 1 to sample noisier versions of $z_{0}$, where $N$ is the number of forward diffusion steps. For each step $n$, we define $\alpha_{n}=1-\beta_{n}$, and calculate the cumulative product $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{n}$.The diffusion equations are described as follows:
@@
where the noise term $\epsilon$ and the final step of the forward process yields $z_{N}$ follow a Gaussian distribution, specifically $\epsilon,z_{N}\sim\mathcal{N}(0,I)$. The reverse process denoises and reconstructs $z_{0}$ through text-guided noise estimation $(\hat{\epsilon}_{\theta})$ using following loss function:
@@
After training LDM, we generate audio latent by sampling through the reverse process with $z_{N}\sim\mathcal{N}(0,I)$, conditioned on the given textual representation $\tau$. Its reverse dynamics are shown below:
@@
After that, the decoder of audio VAE [18] constructs a mel-spectrogram feature from the latent audio representation $\hat{z}_{0}$. This mel-spectrogram feature is conveyed to a vocoder to generate the final audio.
@@ 3.3 Visual-to-Audio Event Understanding Module
The diversity of sound is influenced by various factors such as its source, actions, the environment, and more. At the same time, these factors are often included in the description of visual images. Inspired by the widespread use of VLMs [23, 37, 39], we chose MiniGPT-v2 [2] to process visual information and generate descriptions of sounds. Recognizing that MiniGPT-v2 was initially developed for single-image understanding and its limitations in conveying dynamic information, we have adapted the LLaMA-2 [34] conversation template design to suit a multi-modal instructional framework. To not miss sound events, we encoded four random video frames with time-sensitive prompts (temporal cues):
First ,< Img >< ImageFeature >< Img >.
Then ,< Img >< ImageFeature >< Img >.
After that,< Img >< ImageFeature >< Img >.
Finally ,< Img >< ImageFeature >< Img >.
[Task Identifier ] Instruction
@@ 3.4 Sound Event Timestamp Detection Module
In practice, sound artists need to manually determine the point in time when a sound event starts and ends and then adjust the appropriate sound effect to the correct position. This judgment is usually based on current visual information. To simplify the process, we use the sound event timestamp detection module to detect the timestamp of the sound event in the video inspired by the CondFoleyGen [9]. Instead of using a hand-crafted approach to transfer sounds from the conditional audio, We use a ResNet(2+1)-D18 [35] visual network to capture timestamps as time-conditional inputs to the LDM, trained on paired video and timestamp.
The workflow begins with feeding a sequence of silent video frames $V_{f}$ into the detection network. The network then outputs a binary vector $V_{ct}$ representing predictions for each frame, derived from a fully connected layer post-pooling. The ground truth $V_{ct}$ is obtained: Function $P$ detects audio $a$ in each frame, applying a threshold $x$ to reduce noise effects. Sounds within 0.02s across consecutive frames are considered a single event. This method identifies the start $x_{start}$ and end points $x_{end}$ of sound segments, as delineated in Eq. 7. Based on timestamps, we construct $T_{ct}$, as depicted in Eq. 8, where 1 indicates the presence of sound, and 0 indicates the absence. The process equations are as follows:
$x_{start},x_{end}=P\left(a_{c}\right),$ (7)
$T_{ct}=\begin{cases}1\ \ \ ,if\ \ t\ \ in\ \ \ [x_{start},x_{end}],\\ 0\ \ \ ,else.\end{cases}$ (8)
Finally, $T_{ct}$ is adjusted to correspond with the video frames’ duration, represented as $V_{ct}$. We employ binary cross-entropy loss to penalize inaccuracies in time prediction. Given that an input video may contain multiple sound events, each sound event’s weights are based on its duration relative to the total sound duration in the video. The binary vector $T_{ct}$ serves as input for the Time-controllable Latent Diffusion Model. The whole process and network structure are shown in Fig. 3.
Figure 3: Sound Event Timestamp Detection Module. The network analyzes the video’s features to output a binary vector corresponding to the video’s frame count. Within this vector, sections marked in white (value of 1) mean sound presence, and those in black (value of 0) indicate sound absence.
@@ 3.5 Time-controllable Latent Diffusion Model
In our experimentation, we observed that the results generated were semantically inaccurate and temporally unsynchronized. This issue often arose when utilizing audio-visual datasets to train end-to-end models, limiting the task’s practical application. We attribute these shortcomings to the complexity of the sound sources and the poor audio quality of the audio-visual dataset used for training. To address this, we use text to bridge audio and video and then introduce time control in the T2A generation model. In the visual domain, many works are based on the ControlNet [40] architecture, which can finer control the generation of images or videos by manipulating the input image conditions of neural network blocks. However, unlike the visual domain’s inherent intuitiveness, audio features exhibit temporal continuity and are inherently more complex and abstract. Thus, selecting appropriate audio features to guide the generation process poses a greater challenge.
In this paper, we propose a embedding called Audio Time-condition Embedding, denote as $A_{ct}$, which is constructed through the following procedure: During the training phase, the adapter extracted Mel-spectrogram features from audio waveform $a$ and then normalized it, denoted as $a_{mel}=mel(a)$. For $a_{c}=max(mel(a),T_{ct})$, we then initialize $a_{c}$ to match $a_{mel}$’s dimensions but with zero values, then fill $a_{c}$’s Mel channels with the maximum values based on the corresponding frames where $T_{ct}$ equals 1. Finally, $A_{ct}$ is derived by encoding the embedding of $a_{c}$ via the encoder $E_{a}$, and described as follows:
$A_{ct}=E_{a}\left(\operatorname{max}\left(\operatorname{mel}\left(a_{c}\right),T_{ct}\right)\right).$ (9)
Inspired by ControlNet [40], we have developed a network architecture named Time-controllable Adapter. Then, we integrate the audio time-condition embedding $A_{ct}$ with the text embedding $\tau$ and target audio embedding into the neural network block. This integration facilitates joint training for the time-controllable adapter. $A_{ct}$ is fed solely into the adapter, while text embedding $\tau$ inputs into both Tango and the adapter.
Tango’s denoising model, akin to UNet, includes encoder $F$, middle block $M$, and decoder $G$. $F$ and $G$ have 12 corresponding blocks. Tango’s outputs of the encoder’s $i-th$ block and decoder’s $j-th$ block are $f_{i}$ and $g_{j}$, with $m$ for the single middle block. The adapter mimics Tango’s encoder and middle layers as $F^{\prime}$ and $M^{\prime}$. Adapter’s outputs, like $f_{i}^{\prime}$ and $m^{\prime}$, are marked with (). Following this, the output of the time-controllable adapter is concatenated with the output of the corresponding decoder block during the decoding process. For example, $f_{1}$ from the 1st encoder layer adds to $g_{12}$ of the decoder, making $i+j=13$. To achieve it, we ensure that all Tango’s elements are kept frozen while modifying the input of the $i-th$ block of the decoder as:
$\begin{cases}Concat\left(m+m^{{}^{\prime}},f_{j}+zero\left(f_{j}^{{}^{\prime}}\right)\right)\ \ i=1,j=12.\\ \\ Concat\left(g_{i-1},f_{j}+zero\left(f_{j}^{{}^{\prime}}\right)\right)2\leqslant i\leqslant 12,i+j=13.\\ \end{cases}$ (10)
$zero(\cdot)$ is one zero convolutional layer, facilitating trainable and fixed neural network block connections. Its weights evolve from zero to optimized values during training. This approach not only retains the Tango’s capability for generating audio, trained on billions of audio-text pairs, but also enables the model to comprehend the guidance provided by the time control embedding, resulting in temporally controllable outcomes.
Table 1: Distribution of CondPromptBank Categories
Category % Category % Category %
Household Daily 14.11 Transportation Vehicles 10.42 Impacts Crashes 10.10
Foley 8.24 Human Elements 7.77 Industrial 6.58
Weapons War 5.83 Cartoon Comical 4.90 Sports 4.43
Animals Insects 4.04 Instruments 3.68 Water Liquid 3.27
Technology 2.70 Horror 2.41 Emergency 2.20
Public Places 1.87 Sound Design Effects 1.69 Doors Windows 1.56
Fire Explosions 1.49 Nature Weather 1.02 Leisure 0.84
Multimedia 0.47 Bells 0.37
@@
CondFoleyGen(old) [9] 17.3 15.6 21.6 67.0 35.5 23.3
CondFoleyGen(new) [9] 16.3 13.6 19.2 68.6 37.4 22.4
Ours-small 29.6 28.0 19.4 77.03 27.8 35.6
Ours-full 36.8 42.8 27.6 78.1 43.8 39.7
Table 2: Conditional Generation Task Quantitative Results. CLAP-top metric evaluates a model’s ability to control sound content. It calculates the percentage of times the sound samples generated by all models ranked in the top 1 according to the CLAP ranking, divided by the total number of samples. We used two versions of the CLAP models for evaluation: $CLAP_{general}$ and $CLAP_{unfused}$. Onset Acc and Time Acc are both metrics based on the number of sound occurrences, with Onset Acc focusing on the onset count and Time Acc on the count of time intervals. We measure the average accuracy of predicting onsets within 0.1s of ground truth to assess the timing of generated onsets. IoU is calculated by computing the intersection and union of these vectors. These metrics collectively allow us to comprehensively evaluate the accuracy of the generated sounds in terms of both timing and content. The variant “old" corresponds to the prior codebook, while variant “new" matches the updated codebook. The old model was trained on 192-width spectrograms, and the new one was trained on 2s waveforms. “ ”denotes the data sourced from the official code and is based on experiments conducted with our local configurations. Underline denotes the worst performance. Boldface denotes the best performance. The six metrics are all measured in percentage.